# Kilohertz Gravitational Waves from Binary Neutron Star Mergers: Full Spectrum Analyses and High-density Constraints on Neutron Star Matter

### v1

@@ III.3 EoS constraints
In this section we present the results of the EoS study enabled by our PE analyses. In particular EoS sampling can be performed starting from the results of the IM-only PEs, employing only the masses and the tidal polarizability parameters, or from the results of the PM, incorporating also the constraints from the quasi-universal relations for the maximum density ${\rho_{\rm max}^{\rm TOV}}$ and the radius at the maximum mass ${R_{\rm max}^{\rm TOV}}$, as described in Sec. II.3.
@@
We begin examining the distribution of the maximum density and the radius at the maximum mass, computed from the values of $f_{2}$ using the quasi-universal relations (9) and (10), respectively. Figure 9 displays the posterior distributions of $\Delta{\rho_{\rm max}^{\rm TOV}}=\rho_{\rm max}^{\rm TOV,PE}-{\rho_{\rm max}^{\rm TOV}}$ and $\Delta{R_{\rm max}^{\rm TOV}}=R_{\rm max}^{\rm TOV,PE}-{R_{\rm max}^{\rm TOV}}$, for five different EoSs. ${\rho_{\rm max}^{\rm TOV}}$ and ${R_{\rm max}^{\rm TOV}}$ corresponds to the maximal theoretical values obtained from the stable EoS sequence, while $\rho_{\rm max}^{\rm TOV,PE}$ and $R_{\rm max}^{\rm TOV,PE}$ are computed from the fits. We observe that for soft EoSs the distributions of $\Delta{\rho_{\rm max}^{\rm TOV}}$ are exactly centered around zero, whereas for stiffer EoSs, such as DD2, the distributions are shifted away from zero, although they remain compatible with zero within the $90\%$ CI. This discrepancy arises from the fit relations used in the analysis. In particular examining Fig. 1 of Breschi et al. (2022a), we note that the specific binary with DD2 EoS we used exhibits a larger deviation from the fitted relation for ${\rho_{\rm max}^{\rm TOV}}$, although it remains within the 90% CI.
@@
Combining the posteriors of the total mass and the tidal parameters with the results above, we constrain the mass-radius diagram. Figure 10 shows the $M(R)$ relation for the SLy ($\mathcal{M}=1.244295~{}{\rm M_{\odot}}$) and BLh ($\mathcal{M}=1.227164~{}{\rm M_{\odot}}$) EoSs (See Table 2). The constrained areas are obtained with TEOBResumSPA for IM-only and TEOBResumSPA_NRPMw for IMPM. For PM-only we use the TEOBResumSPA_NRPMw PE to compute ${R_{\rm max}^{\rm TOV}}$ and ${\rho_{\rm max}^{\rm TOV}}$, but we are not including the constraints on the mass and the tidal polarizability parameter. We utilize the EoSs prior sample of Godzieba et al. (2021) and in our 90% CI 14372 EoSs are included. We additionally test scenarios varying this amount, i.e. by doubling it, and obtain consistent results.
@@
The PM gives information at higher densities and together with the IM produces a tighter constraint around the theoretical mass-radius curve, giving in particular a better estimation of the maximum mass ${M_{\rm max}^{\rm TOV}}$ and the corresponding radius ${R_{\rm max}^{\rm TOV}}$. For IMPM, we find the CIs to be ${M_{\rm max}^{\rm TOV}}=2.03^{+0.07}_{-0.05}~{}{\rm M_{\odot}}$ and ${R_{\rm max}^{\rm TOV}}=10.23^{+0.25}_{-0.12}{\rm km}$ for SLy and ${M_{\rm max}^{\rm TOV}}=2.07^{+0.09}_{-0.08}~{}{\rm M_{\odot}}$ and ${R_{\rm max}^{\rm TOV}}=10.75^{+0.32}_{-0.25}~{}{\rm km}$ for BLh. The values we obtain are compatible with Breschi et al. (2021) and the relative error is roughly the same. Compared with earlier analyses based on GW170817 Breschi et al. (2024b); Nathanail et al. (2021); Koehn et al. (2025); Essick et al. (2020), an improvement by a factor three in the uncertainties of the maximum mass is achieved with XG detectors. For ${R_{\rm max}^{\rm TOV}}$, the precision enhances by a factor four Rutherford et al. (2024); Fan et al. (2024). We note that in the region of $1.54{\rm M_{\odot}}\lesssim M\lesssim 1.87{\rm M_{\odot}}$, the constrained IMPM 90% confidence region is not totally consistent with the theoretical mass-radius relation for SLy and BLh. We attribute this slight incompatibility to the sample of EoSs, that does not include EoSs exactly compatible with SLy or BLh at $M\lesssim 0.6~{}{\rm M_{\odot}}$.
@@
Furthermore, Figure 11 reports the $\Lambda(M)$ curve derived from the combined IMPM analysis, evaluated for an equal-mass binary with BLh EoS, $M=2.8{\rm M_{\odot}}$ and $\tilde{\Lambda}=412$. The resulting constraint on the tidal polarizability parameter as a function of the mass is particularly tight around the theoretical curve for BLh, highlighting both the accuracy of our method in recovering the theoretical prediction and its robustness in probing the dense matter properties of NSs. From this analysis, we also extract constraints on the tidal polarizability parameter at $M=1.4{\rm M_{\odot}}$, finding that it lies within the range $529289<\Lambda_{1.4}<$ 529.
@@
Finally, Figure 12 displays an example of direct inference of the pressure-density EoS relation. The theoretical SLy curve lies within the 90% CIs for both IM and IMPM cases. We observe a similar behavior to that seen in the mass-radius curve: the confidence region is narrower in the region with $5\times 10^{14}~{}{\rm g~{}cm^{-3}}\lesssim\rho\lesssim 13\times 10^{14}~{}{\rm g~{}cm^{-3}}$, while it becomes broader at both lower and higher densities. As in the other inferences discussed, the IMPM analysis improves the precision of the inference on $p(\rho)$ at higher densities.
@@ IV Conclusion
Using an EoS sampling with minimal hypotheses (causality, validity of general relativity), we showed that a single-event PM signal plays a crucial role in constraining both the maximum density and the radius corresponding to the maximum mass, through quasi-universal relations that involves the peak frequency $f_{2}$. The inclusion of these posterior distributions significantly reduces the 90% credibility region of the mass-radius diagram, particularly at high-densities. Therefore, the possibility to perform PEs of GW signals detected with ET, that also contains PM features, enables more precise constraints on the maximum mass and the corresponding radius. In our study, we demonstrate relative errors smaller than 8% and 5% for the maximum mass and the relative neutron star radius, respectively. A novel result presented in this paper is the tight constraint on the $\Lambda(M)$ curve, where the 90% CI closely follows the theoretical model. Similarly, the inference of the pressure-density relation, especially that obtained through the IMPM analysis, yields significantly tight constraints at high-densities, again closely tracking the theoretical prediction. This approach enhances the precision in extracting information about the properties of nuclear matter under extreme conditions.
@@
Our work has important limitations that should be considered for future work. First, in both IM-only and full spectrum analyses, we injected non-spinning binary systems. In the future, it would be valuable to access the ability of our pipeline to accurately recover spin components, despite the degeneracy between spin, mass ratio and tidal polarizability parameters. An ongoing study is already investigating PE without fixing any parameters in the IMPM scenario. This includes the inference of the spin components as well as additional extrinsic parameters of the binary, beyond just the luminosity distance. Moreover, our current study focuses only on the dominant (2,2) mode. Including also higher-order modes would help, for example to address the degeneracy between luminosity distance and the inclination. TEOBResumSPA already incorporates higher modes, whereas their implementation in NRPMw is still undergoing. Given that our injections are based on GW signals from XG of detectors, we must also account for the long duration of such signal. In this regime, Earth’s rotation is not negligible, and the increased sensitivity of the detector implies that multiple signals may overlap in the data stream. Consequently, future analysis techniques must address challenges in signal disentanglement and noise mitigation Abac et al. (2025); Cireddu et al. (2024); Liu et al. (2025); Wong et al. (2025). Lastly, in our analyses, we used the same model for both injection and recovery. However, Gamba et al. (2021b) showed that for signals with SNR $>80$, using different waveform models for injection and recovery introduces significant systematic errors. This highlights the importance of waveform systematics studies and motivates the development of more robust cross-model validation strategies.
@@ Acknowledgments
The authors thank Nishkal Rao for technical contributions and discussion. GH and SB acknowledge support by the EU Horizon under ERC Consolidator Grant, no. InspiReM-101043372. SB and MB acknowledge support by the EU H2020 under ERC Starting Grant, no. BinGraSp-714626. SB acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) project MEMI (BE 6301/2-1 Projektnummer: 443239082) and GROOVHY (BE 6301/5-1 Projektnummer: 523180871). MB and RG acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) under Grant No. 406116891 within the Research Training Group RTG 2522/1. RG acknowledges support from NSF Grant PHY-2020275 (Network for Neutrinos, Nuclear Astrophysics, and Symmetries (N3AS)).
Computations were performed on SuperMUC-NG at the Leibniz-Rechenzentrum (LRZ) Munich and and on the national HPE Apollo Hawk at the High Performance Computing Center Stuttgart (HLRS). The authors acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at LRZ (allocations pn76li, pn36jo and pn68wi). The authors acknowledge HLRS for funding this project by providing access to the supercomputer HPE Apollo Hawk under the grant number INTRHYGUE/44215 and MAGNETIST/44288. Postprocessing and development runs were performed on the ARA cluster at Friedrich Schiller University Jena. The ARA cluster is funded in part by DFG grants INST 275/334-1 FUGG and INST 275/363-1 FUGG, and ERC Starting Grant, grant agreement no. BinGraSp-714626.
@@ Appendix A Full spectrum model
Long-lived SLy Breschi et al. (2019) $2.6$ $1.0$ $92$ $3.13$ $21$
@@
Figure 14: Unfaithfulness between hybrid NR-EOB waveforms and TEOBResumSPA_NRPMw model as function of the lower cut-off frequency $f_{\rm min}$. The values are computed employing ET-D Hild et al. (2011); Hild (2012) sensitivity curve and setting the upper cut-off frequency to $8~{}{\rm kHz}$. The vertical dotted lines refer to the value of the instantaneous NR frequency at merger for each case.
@@
The validation of the complete model has been performed using 6 exemplary NR simulations from the CoRe database Gonzalez et al. (2023) computed with the thc code Bernuzzi et al. (2016); Perego et al. (2019); Bernuzzi et al. (2020); Prakash et al. (2021); Kashyap et al. (2022) and listed in Tab. 3. Note this set of simulations is not included in the calibration set. The data are chosen to cover a sufficiently large variety of phenomenologies; in particular, the validation set includes four long-lived cases, two of which with $q>1.5$ showing tidal disruption, and two short-lived cases, one of which includes deconfined quark matter. Moreover, NR data are hybridized with the time-domain TEOBResumS model, in order to extended the waveforms in the early-inspiral regime, not included in typical NR simulations. Also for these studies, we employ ET-D sensitivity curve Hild et al. (2011); Hild (2012) and we minimize over the additional set of parameters $\{{\boldsymbol{\theta}_{\rm PM}^{\rm free}},\boldsymbol{\delta}_{\rm PM}\}$ for NRPMw. Employing only the PM portion of data and NRPMw, we obtain $\log_{10}{\bar{\mathcal{F}}}$ between $-1.35$ and $-0.95$ over all the validation set of 6 binaries. In particular, short-lived cases (SFHo and BLQ) reach $\log_{10}{\bar{\mathcal{F}}}\lesssim-1.25$; while, large mass ratio binaries (BLh and DD2) recover worst matches, i.e.  $\log_{10}{\bar{\mathcal{F}}}\simeq-0.95$.
Figure 14 shows the unfaithfulness computed between TEOBResumSPA_NRPMw model and the hybrid NR-EOB waveforms, where the inner products are evaluated over the frequency range $[f_{\rm min},8~{}{\rm kHz}]$. For $f_{\rm min}<100~{}{\rm Hz}$, the model have good agreement with NR data as shown by $\log_{10}{\bar{\mathcal{F}}}\lesssim-3$, comparable to Ref. Gamba et al. (2021a). Approaching $f_{\rm mrg}$, the recovered ${\bar{\mathcal{F}}}$ shows an increasing trend, with peaks above $10^{-1}$. Subsequently, for $f_{\rm min}>1.5~{}{\rm kHz}$, the recovered ${\bar{\mathcal{F}}}$ are comparable with the values obtained only with the PM portion of data, i.e.  $\log_{10}{\bar{\mathcal{F}}}\lesssim-1$. In order to further inspect this behavior, we show in Fig. 15 the comparison between the time-domain NR data and the corresponding TEOBResumSPA_NRPMw templates, making use of the inverse FFT. The inspiral EOB template matches NR data accurately up to merger, i.e.  $t=0$; however, considerable phase differences arise after this moment in the two models. These errors can be related to the stationary-phase approximation (SPA) and to the absence of next-to-quasi-circular (NQC) corrections Riemenschneider et al. (2021); Albertini et al. (2022) in the frequency-domain EOB template. These corrections appear to be more relevant for equal-mass binaries and they motivate additional studies for the development and the calibration of accurate EOB-NR waveform models.

### v2

@@ III.3 EoS constraints
In this section we present the results of the EoS study enabled by our PE analyses. In particular EoS sampling can be performed starting from the results of the IM-only PEs, employing only the masses and the tidal polarizability parameters, or from the results of the PM, incorporating also the constraints from the quasiuniversal relations for the maximum density $\rho_{\rm max}^{\rm TOV}$} and the radius at the maximum mass $R_{\rm max}^{\rm TOV}$}, as described in Sec. II.3.
@@
We begin examining the distribution of the maximum density and the radius at the maximum mass, computed from the values of $f_{2}$ using the quasiuniversal relations (9) and (10), respectively. Figure 9 displays the posterior distributions of $\Delta{\rho_{\rm max}^{\rm TOV}}=\rho_{\rm max}^{\rm TOV,PE}-{\rho_{\rm max}^{\rm TOV}}$ and $\Delta{R_{\rm max}^{\rm TOV}}=R_{\rm max}^{\rm TOV,PE}-{R_{\rm max}^{\rm TOV}}$, for five different EoSs. $\rho_{\rm max}^{\rm TOV}$} and $R_{\rm max}^{\rm TOV}$} corresponds to the maximal theoretical values obtained from the stable EoS sequence, while $\rho_{\rm max}^{\rm TOV,PE}$ and $R_{\rm max}^{\rm TOV,PE}$ are computed from the fits. We observe that for soft EoSs the distributions of $\Delta{\rho_{\rm max}^{\rm TOV}}$ are exactly centered around zero, whereas for stiffer EoSs, such as DD2, the distributions are shifted away from zero, although they remain compatible with zero within the $90\%$ CL. This discrepancy arises from the fit relations used in the analysis. In particular examining Fig. 1 of Breschi et al. (2022), we note that the specific binary with DD2 EoS we used exhibits a larger deviation from the fitted relation for $\rho_{\rm max}^{\rm TOV}$}, although it remains within the 90% CL.
@@
Combining the posteriors of the total mass and the tidal parameters with the results above, we constrain the mass-radius diagram. Figure 10 shows the $M(R)$ relation for the SLy ($\mathcal{M}=1.244295~{\rm M_{\odot}}$) and BLh ($\mathcal{M}=1.227164~{\rm M_{\odot}}$) EoSs (See Table 2). The constrained areas are obtained with TEOBResumSPA for IM-only and TEOBResumSPA_NRPMw for IMPM. For PM-only we use the TEOBResumSPA_NRPMw PE to compute $R_{\rm max}^{\rm TOV}$} and $\rho_{\rm max}^{\rm TOV}$}, but we are not including the constraints on the mass and the tidal polarizability parameter. We utilize the EoSs prior sample of Godzieba et al. (2021) and in our 90% CL 14372 EoSs are included. We additionally test scenarios varying this amount, i.e. by doubling it, and obtain consistent results.
@@
Comparing the results for IM-only, PM-only and IMPM of Fig. 10 we can observe the impact of the constraints given by the TEOBResumSPA_NRPMw waveform model alone and by adding the quasiuniversal relations. The IM-only case is based only on the results from the GW analysis, while the PM-only is based solely on the inference on $f_{2}$ and the quasiuniversal relations. The IMPM case combines the two in a stricter constraint. Therefore going from the case of the IM-only (yellow) to the IMPM (red) we additionally include the constraints coming from the quasiuniversal relation to the GW inference of $M$ and $\tilde{\Lambda}$. Looking at Fig. 7, we can see that the inference of masses and tidal polarizability parameters is dominated by the IM. The case of the full-spectrum model TEOBResumSPA_NRPMw does not give additional constraints. Had we repeated the EoS sampling with the IMPM without any information from the quasiuniversal relations from $f_{2}$, we would have obtained constraints identical to the IM-only. The PM, instead, gives information at higher densities and together with the IM produces a tighter constraint around the theoretical mass-radius curve, giving in particular a better estimation of the maximum mass $M_{\rm max}^{\rm TOV}$} and the corresponding radius $R_{\rm max}^{\rm TOV}$}. For IMPM, we find the CLs to be ${M_{\rm max}^{\rm TOV}}=2.03^{+0.07}_{-0.05}~{\rm M_{\odot}}$ and ${R_{\rm max}^{\rm TOV}}=10.23^{+0.25}_{-0.12}{\rm km}$ for SLy and ${M_{\rm max}^{\rm TOV}}=2.07^{+0.09}_{-0.08}~{\rm M_{\odot}}$ and ${R_{\rm max}^{\rm TOV}}=10.75^{+0.32}_{-0.25}~{\rm km}$ for BLh. The values we obtain are compatible with Breschi et al. (2021) and the relative error is roughly the same. Compared with earlier analyses based on GW170817 Breschi et al. (2024b); Nathanail et al. (2021); Koehn et al. (2025); Essick et al. (2020), an improvement by a factor three in the uncertainties of the maximum mass is achieved with XG detectors. For $R_{\rm max}^{\rm TOV}$}, the precision enhances by a factor four Rutherford et al. (2024); Fan et al. (2024). We note that in the region of $1.54{\rm M_{\odot}}\lesssim M\lesssim 1.87{\rm M_{\odot}}$, the constrained IMPM 90% confidence region is not totally consistent with the theoretical mass-radius relation for SLy and BLh. We attribute this slight incompatibility to the sample of EoSs, that does not include EoSs exactly compatible with SLy or BLh at $M\lesssim 0.6~{\rm M_{\odot}}$.
@@
Furthermore, Fig. 11 reports the $\Lambda(M)$ curve derived from the combined IMPM analysis, evaluated for an equal-mass binary with BLh EoS, $M=2.8{\rm M_{\odot}}$ and $\tilde{\Lambda}=412$. The resulting constraint on the tidal polarizability parameter as a function of the mass is particularly tight around the theoretical curve for BLh, highlighting both the accuracy of our method in recovering the theoretical prediction and its robustness in probing the dense matter properties of NSs. From this analysis, we also extract constraints on the tidal polarizability parameter at $M=1.4{\rm M_{\odot}}$, finding that it lies within the range $289<\Lambda_{1.4}<529$.
@@
Finally, Fig. 12 displays an example of direct inference of the pressure-density EoS relation. The theoretical SLy curve lies within the 90% CLs for both IM and IMPM cases. We observe a similar behavior to that seen in the mass-radius curve: the confidence region is narrower in the region with $5\times 10^{14}~{\rm g~cm^{-3}}\lesssim\rho\lesssim 13\times 10^{14}~{\rm g~cm^{-3}}$, while it becomes broader at both lower and higher densities. As in the other inferences discussed, the IMPM analysis improves the precision of the inference on $p(\rho)$ at higher densities.
@@ IV Conclusion
Using an EoS sampling with minimal hypotheses (causality, validity of general relativity), we showed that a single-event PM signal plays a crucial role in constraining both the maximum density and the radius corresponding to the maximum mass, through quasiuniversal relations that involves the peak frequency $f_{2}$. The inclusion of these posterior distributions significantly reduces the 90% credibility region of the mass-radius diagram, particularly at high densities. Therefore, the possibility to perform PEs of GW signals detected with ET, that also contains PM features, enables more precise constraints on the maximum mass and the corresponding radius. In our study, we demonstrate relative errors smaller than 8% and 5% for the maximum mass and the relative neutron star radius, respectively. A novel result presented in this paper is the tight constraint on the $\Lambda(M)$ curve, where the 90% CL closely follows the theoretical model. Similarly, the inference of the pressure-density relation, especially that obtained through the IMPM analysis, yields significantly tight constraints at high densities, again closely tracking the theoretical prediction. This approach enhances the precision in extracting information about the properties of nuclear matter under extreme conditions.
@@
Our work has important limitations that should be considered for future work. First, in both IM-only and full spectrum analyses, we injected nonspinning binary systems. In the future, it would be valuable to access the ability of our pipeline to accurately recover spin components, despite the degeneracy between spin, mass ratio and tidal polarizability parameters. An ongoing study is already investigating PE without fixing any parameters in the IMPM scenario. This includes the inference of the spin components as well as additional extrinsic parameters of the binary, beyond just the luminosity distance. Moreover, our current study focuses only on the dominant (2,2) mode. Including also higher-order modes would help, for example to address the degeneracy between luminosity distance and the inclination. TEOBResumSPA already incorporates higher modes, whereas their implementation in NRPMw is still undergoing. Given that our injections are based on GW signals from XG of detectors, we must also account for the long duration of such signal. In this regime, Earth’s rotation is not negligible, and the increased sensitivity of the detector implies that multiple signals may overlap in the data stream. Consequently, future analysis techniques must address challenges in signal disentanglement and noise mitigation Abac et al. (2025); Cireddu et al. (2024); Liu et al. (2025); Wong et al. (2025). Lastly, in our analyses, we used the same model for both injection and recovery. However, Gamba et al. (2021b) showed that for signals with SNR >80>80, using different waveform models for injection and recovery introduces significant systematic errors. This highlights the importance of waveform systematics studies and motivates the development of more robust cross-model validation strategies.
@@ Acknowledgments
The authors thank Nishkal Rao for technical contributions and discussion. G.H. and S.B. acknowledge support by the EU Horizon under ERC Consolidator Grant, No. InspiReM-101043372. S.B. and M.B. acknowledge support by the EU H2020 under ERC Starting Grant, no. BinGraSp-714626. S.B. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) project MEMI (BE 6301/2-1 Projektnummer: 443239082) and GROOVHY (BE 6301/5-1 Projektnummer: 523180871). M.B. and R.G. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) under Grant No. 406116891 within the Research Training Group RTG 2522/1. R.G. acknowledges support from NSF Grant No. PHY-2020275 [Network for Neutrinos, Nuclear Astrophysics, and Symmetries (N3AS)].
Computations were performed on SuperMUC-NG at the Leibniz-Rechenzentrum (LRZ) Munich and and on the national HPE Apollo Hawk at the High Performance Computing Center Stuttgart (HLRS). The authors acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at LRZ (allocations pn76li, pn36jo and pn68wi). The authors acknowledge HLRS for funding this project by providing access to the supercomputer HPE Apollo Hawk under the grant number INTRHYGUE/44215 and MAGNETIST/44288. Postprocessing and development runs were performed on the ARA cluster at Friedrich Schiller University Jena. The ARA cluster is funded in part by DFG grants No. INST 275/334-1 FUGG and No. INST 275/363-1 FUGG, and ERC Starting Grant, grant agreement no. BinGraSp-714626.
@@ Appendix A Full spectrum model
Long-lived SLy Breschi et al. (2019) $2.6$ $1.0$ $92$ $3.13$ 2121
@@
Figure 14: Unfaithfulness between hybrid NR-EOB waveforms and TEOBResumSPA_NRPMw model as function of the lower cutoff frequency $f_{\rm min}$. The values are computed employing ET-D Hild et al. (2011); Hild (2012) sensitivity curve and setting the upper cutoff frequency to $8~{\rm kHz}$. The vertical dotted lines refer to the value of the instantaneous NR frequency at merger for each case.
@@
The validation of the complete model has been performed using 6 exemplary NR simulations from the CoRe database Gonzalez et al. (2023) computed with the thc code Bernuzzi et al. (2016); Perego et al. (2019); Bernuzzi et al. (2020); Prakash et al. (2021); Kashyap et al. (2022) and listed in Tab. 3. Note this set of simulations is not included in the calibration set. The data are chosen to cover a sufficiently large variety of phenomenologies; in particular, the validation set includes four long-lived cases, two of which with $q>1.5$ showing tidal disruption, and two short-lived cases, one of which includes deconfined quark matter. Moreover, NR data are hybridized with the time-domain TEOBResumS model, in order to extended the waveforms in the early inspiral regime, not included in typical NR simulations. Also for these studies, we employ ET-D sensitivity curve Hild et al. (2011); Hild (2012) and we minimize over the additional set of parameters $\{{\boldsymbol{\theta}_{\rm PM}^{\rm free}},\boldsymbol{\delta}_{\rm PM}\}$ for NRPMw. Employing only the PM portion of data and NRPMw, we obtain $\log_{10}{\bar{\mathcal{F}}}$ between 1.35-1.35 and 0.95-0.95 over all the validation set of 6 binaries. In particular, short-lived cases (SFHo and BLQ) reach $\log_{10}{\bar{\mathcal{F}}}\lesssim-1.25$; while, large mass ratio binaries (BLh and DD2) recover worst matches, i.e.  $\log_{10}{\bar{\mathcal{F}}}\simeq-0.95$.
Figure 14 shows the unfaithfulness computed between TEOBResumSPA_NRPMw model and the hybrid NR-EOB waveforms, where the inner products are evaluated over the frequency range $f_{\rm min},8~{\rm kHz}]$. For $f_{\rm min}<100~{\rm Hz}$, the model have good agreement with NR data as shown by $\log_{10}{\bar{\mathcal{F}}}\lesssim-3$, comparable to Ref. Gamba et al. (2021a). Approaching $f_{\rm mrg}$, the recovered $\bar{\mathcal{F}}$} shows an increasing trend, with peaks above $10^{-1}$. Subsequently, for $f_{\rm min}>1.5~{\rm kHz}$, the recovered $\bar{\mathcal{F}}$} are comparable with the values obtained only with the PM portion of data, i.e.  $\log_{10}{\bar{\mathcal{F}}}\lesssim-1$. In order to further inspect this behavior, we show in Fig. 15 the comparison between the time-domain NR data and the corresponding TEOBResumSPA_NRPMw templates, making use of the inverse FFT. The inspiral EOB template matches NR data accurately up to merger, i.e.  $t=0$; however, considerable phase differences arise after this moment in the two models. These errors can be related to the stationary-phase approximation (SPA) and to the absence of next-to-quasicircular (NQC) corrections Riemenschneider et al. (2021); Albertini et al. (2022) in the frequency-domain EOB template. These corrections appear to be more relevant for equal-mass binaries and they motivate additional studies for the development and the calibration of accurate EOB-NR waveform models.